OpenStudy (anonymous):

alg 2 help? pleasee. ill fan & medal you :)

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Create two additional quadratic functions, g(x) and h(x). The function g(x) will open the same direction as f(x), have the same vertex, but will be narrower. The function h(x) will open in the opposite direction as f(x), have the same vertex, but will be wider than f(x). Write your functions below and explain in complete sentences why those functions will meet the requirements in the question. Graph your functions. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Explain the appropriate translations of your graphs, if you were born two months earlier and 5 days later. Use complete sentences.

OpenStudy (anonymous):

my birthday is 02 09 :))

OpenStudy (anonymous):

@jojo7 @math&ing001 are you guys going to help me?

OpenStudy (math&ing001):

I'll see what I can do.

OpenStudy (anonymous):

thank youu<3333

OpenStudy (math&ing001):

Do you know what the vertex form is ?

OpenStudy (lovelyanna):

Can you help me @math&ing001 when you are done here?

OpenStudy (math&ing001):

Ok

OpenStudy (lovelyanna):

<------ My question is the third on the list

OpenStudy (anonymous):

No i dont :/ lol.

OpenStudy (math&ing001):

@kendallnicole this is the vertex form of a parabola y=a(x−h)^2+k Now all you need to do is replace h will be your birth month and k with your birth day. You can choose a randomly between 4 and -4.

OpenStudy (anonymous):

so that is for number 1?

OpenStudy (math&ing001):

Yeah !

OpenStudy (anonymous):

thanks!:) number 2?

OpenStudy (math&ing001):

Distribute the function you got before to get it to standard form.

OpenStudy (anonymous):

can you show me?:(

OpenStudy (math&ing001):

Ok show me what you got for the function.

OpenStudy (anonymous):

y=1(x−2)^2+9

OpenStudy (anonymous):

that?

OpenStudy (math&ing001):

Yeah ! Now use this formula (a-b)^2=a^2-2ab+b^2 to get it to standard form.

OpenStudy (anonymous):

Thanks<333333

OpenStudy (math&ing001):

yw

OpenStudy (anonymous):

can you explain how to get it back in standard form? like i understand it, but i have to say it in complete sentences.

OpenStudy (math&ing001):

To convert equation to standard form we simply expand and simplify the binomial square.

OpenStudy (anonymous):

Thanks!!! can you help me with #3?

OpenStudy (math&ing001):

Sure ! For g(x) to be narrower than f(x), you need to choose a bigger 'a'. For h(x) to be wider keep the same vertex and choose 'a' smaller but positive like (0.5 or 0.25...)

OpenStudy (anonymous):

THANK YOU YOURE SAVING MY LIFEEEE

OpenStudy (anonymous):

can you show me how to graphit?!

OpenStudy (anonymous):

@kagıtucak

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Vertex form makes it easier to graph a parabola